Operating Systems and Systems Integration

Quiz 10: Secure Shell

Do not refer to any printed material. No talking or discussion is allowed until the answer sheets are all collected.

Review:

1. Only one of the following regular expressions matches a line that contains exactly five characters:

(A) '.....'
(B) '{5}'
(C) '*****'
(D) '?????'
(E) '^.....$'

2. Which of the following regular expressions matches at least three letter ‘a’s anywhere on the line, i.e., even if there are any number of other characters between the ‘a’s?

(A) 'a.*a.*a'
(B) 'aaa'
(C) 'a{3}'
(D) 'a?a?a'
(E) '^AAA$'

3. awk is:

(A) a programming language useful for separating input data into columns and operating on them (B) short for \underline{awk}ward, because the syntax is very different from Java, and is rather awkward to

learn (C) short for “\underline{a}pples, \underline{w}atermellon and \underline{k}umquat” (D) does not support regular expressions (E) another word for metacharacter

Preparation:

4. Secure Shell is:

(A) what Nick uses to connect to his desktop machine to show the lectures and laboratory work (B) not useful for remote administration (C) always requires you to use a special hardware password device to authenticate yourself, rather

like a pocket calculator (D) is just another name for bash (E) is not useful for port forwarding, or carrying other network traffic, and does not use encryption

5. If you cannot connect to a remote computer using secure shell (ssh), you should:

(A) Call the police (B) Check whether the secure shell server is running on the remote computer, because an ssh

client needs a server to connect to (C) Shut down your own computer immediately and check for security violations (D) Check whether your computer has aids (Acquired Immune Deficiency Syndrome) (E) Use the \underline{S}uper \underline{S}oftware \underline{H}ickup remover to get your SSH working properly again.

Nick Urbanik <nicku(at)nicku.org>  ver. 1.0